OpenStudy (anonymous):

What is the geometric mean of 3 and 48?

OpenStudy (lalaly):

12

OpenStudy (anonymous):

how

OpenStudy (anonymous):

\[\sqrt{3\times 48}\]

OpenStudy (anonymous):

and \[\sqrt{3\times 48}=\sqrt{144}=12\] that is how lalay got it

OpenStudy (anonymous):

oh ok thanks so that is the right answer?

OpenStudy (anonymous):

yes it is right for sure

OpenStudy (anonymous):

ok great could you also post on my other question how to find t9 for the geometric sequence